OpenStudy (anonymous):

Suppose you invest $2500 at an annual interest rate of 3% compounded continously. How much will you have in the account after 7 years.

OpenStudy (amistre64):

Pert should work fine

OpenStudy (amistre64):

\[A=Pe^{rt}\] \[A=2500~e^{.03*7}\]

OpenStudy (texaschic101):

here is an example if you want to check it out... http://cs.selu.edu/~rbyrd/math/continuous/

OpenStudy (anonymous):

I got $3084.20

OpenStudy (ivettef365):

that is correct

OpenStudy (anonymous):

A=2500(2.718)^0.07(7) A=2500(2.718)^0.21 A=2500(1.23367806) A=3084.20?

OpenStudy (anonymous):

thanks

OpenStudy (ivettef365):

should be .03(7)

OpenStudy (anonymous):

yeah you are right

OpenStudy (ivettef365):

:)